# Nature of Andreev bound states in Josephson junctions of triple-point semimetals

### v1

@@ II.2 Results
The thin-barrier limit is physically equivalent to a Dirac-delta potential barrier. However, since we do not have any constraint on the derivatives of the componets of the wavefunction across the junctions, which follows from the linear-in-momentum dispersion, the standard delta-function-potential approximation [43, 47, 40] for thin barriers cannot be used here [36]. Instead, we need to start with Eq. (II.1), and impose the approximations
@@
From the four combinations of the components of the wavefunction appearing at each of the two boundaries, as shown in Eq. (II.1), we get $2\times 4=8$ linear homogeneous equations in the $8$ variables $\left\{a_{l},\,b_{l},\,a,\,b,\,c,\,d,\,a_{r},\,b_{r}\right\}$, which constitute the $8$ unknown coefficients of the wavefunction. The overall $x$-independent factors of $e^{i\,k_{y}\,y}$ cancel out equation by equation. Let $M$ be the $8\times 8$ matrix constructed from the coefficients of these 8 coefficients, whose explicit form is given by
@@
The consistency of the 8 equations is ensured by the condition $\text{det}\,M=0$. After some row and column operations on the matrix $M$, we can bring it to the form $\text{det}\,\tilde{M}=0$, where
@@
$\mathcal{B}=\sin^{2}\chi\left(\cos^{4}\theta_{n}+\cos^{4}\theta_{r}\right)-\left[3\cos(2\chi)+1\right]\cos^{2}\theta_{n}\cos^{2}\theta_{r}$
@@
In Fig. 2, we show the variation of $|\varepsilon|$ as a function of (a) $k_{\perp}$ (with $\varphi_{12}$ held fixed), and (b) $\phi_{12}$ (with $k_{\perp}$ held fixed), for some representative values of $V_{0}$ and $L$. Since $\mathcal{C}$ is a function of $\cos\varphi_{12}$ [cf. Eq. (II.2)], it follows from Eq. (28) that the subgap energies are periodic in $\varphi$ with period $2\pi$, which is also evident from the curves shown in the lower panel of Fig. 2. In addition, we have plotted the behaviour of ±$\pm|\varepsilon|$ as functions of $\varphi$ and $k_{\perp}$ in the upper panel of Fig. 3, thus illustrating the dependence of the subgap energies on both these variables in a combined way. Subfigures Fig. 3(a) and Fig. 3(b) show the behaviour for two sets of values of the remaining junction variables $V_{0}$ and $L$, with $E_{F}$ kept fixed at unity.
@@
where $\epsilon_{n}$ labels the energy values of the four ABSs, arising from the two distinct values of $|\varepsilon|$. Here, $f(\zeta)=1/\left(1+e^{\frac{\zeta}{k_{B}\,T}}\right)$ is the Fermi-Dirac distribution function, with $k_{B}$ being the Boltzmann constant. The lower panel of Fig. 3 shows the behaviour of $I_{J}$ as a function of $\varphi$, scaled by the appropriate numbers/variables (which we label as $\tilde{I}$), for two sets of representative parameters.
@@
The 3d pseudospin-1 quasiparticles are the generalizations of the pseudospin-1/2 quasiparticles in Weyl semimetals, with the nodal points acting as sources/sinks of Berry flux, acting as Berry curvature monopoles of magnitude 2 [3, 6]. They can be realized in 3d tight-binding models for cold fermionic atoms in cubic optical lattices [9] and, via ab initio simulations, have been identified in materials like TaN, NbN, and WC-type ZrTe [6, 7, 13, 8]. Analogous to Eq. (1), the effective low-energy continuum Hamiltonian, in the vicinity of a nodal point, is given by
@@
The 3d current density operator takes the form $\check{\mathbf{j}}_{3d}(\mathbf{k})=\nabla_{\mathbf{k}}\mathcal{H}_{3d}(\mathbf{k})=\mathcal{S}_{x}\,\hat{\mathbf{x}}+\mathcal{S}_{y}\,\hat{\mathbf{y}}+\mathcal{S}_{z}\,\hat{\mathbf{z}}$. The current density vector from the wavefunction $\psi\equiv\left(c_{1}\quad c_{2}\quad c_{3}\right)^{T}$ is
@@
This again leads to a vanishing contribution from the flat-band wavefunction.
Eq. (32) also tells us that if the quasiparticles encounter a perpendicular barrier, say at $z=z_{0}$, while propagating along the $z$-direction, the conservation of the probability current density leads to the two boundary conditions [20]
$\lim_{\delta\rightarrow 0}c_{1}(x_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{1}(x_{0}+\delta)\text{ and }\lim_{\delta\rightarrow 0}c_{3}(x_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{3}(x_{0}+\delta)\,,$ (33)
which we will implement in our derivations. In this case, if the Fermi energy cuts the dispersion profile at an energy value $E$, the wavefunction will have the exponential factor $e^{i\,\text{sgn}(E)\,k_{z}\,z}$, where $k_{z}=\sqrt{E^{2}-k_{\perp}^{2}}$ and $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$.
III.1 S-B-S junctions oriented perpendicular to the $z$-direction
Here we consider the propagation of the quasiparticles/quasiholes along the $z$-axis. Setting up the S-B-S configuration as shown in Fig. 1(b), we model the superconducting pair potential as
@@ III.1 S-B-S junctions oriented perpendicular to the $z$-direction
where the subscripts $\{1,2,3\}$ on the fermionic creation and annihilation operators represent the three distinct band indices. Analogous to the 3d case, we assume the energy-scale hierarchies $V_{0}\gg E_{F}\gg\Delta_{0}$ and $(V_{0}-E_{F})\gg E_{F}$.
@@
expressed in a piecewise manner for the three regions, where we set the the Fermi energy at $E_{F}$ for the corresponding normal states (i.e., for $\Delta_{0}=0$) in the regions I and III.
@@
Since the final results depend on the phase difference $\varphi_{12}=\varphi_{2}-\varphi_{1}$, for simplification of the notations, we can set $\varphi_{1}=0$ and $\varphi_{2}=\varphi_{12}$, without any loss of generality. Imposing the continuity of the probability current density at the junctions located at $z=0$ and $z=L$, we get the following conditions [cf. Eq. (33)]:
@@ III.2 Results
In order to implement the thin-barrier limit, we start with Eq. (6), and impose the approximations
@@
From the four combinations of the components of the wavefunction appearing at each of the two boundaries, as shown in Eq. (III.1), we get $2\times 4=8$ linear homogeneous equations in the $8$ variables $\left\{a_{l},\,b_{l},\,a,\,b,\,c,\,d,\,a_{r},\,b_{r}\right\}$, which constitute the $8$ unknown coefficients of the wavefunction. In the resulting equations, both the overall $z$-independent factors of $e^{i\left(k_{x}\,x\,+\,k_{y}\,y\right)}$ and the phase factors introduced by the action of $e^{-i\,\phi\,{\mathcal{S}}_{0}\otimes{\mathcal{S}}_{z}}$ cancel out component by component. Let $\check{M}$ be the $8\times 8$ matrix constructed from the coefficients of these 8 coefficients, whose explicit form is given by
@@
The consistency of the 8 equations is ensured by the condition $\text{det}\,\check{M}=0$. After some row and column operations on the matrix $M$, we can bring it to the form which turns out to be the same as Eq. (26). Hence, the results for the 2d case ABSs are valid for the 3d case as well.
@@ IV Summary and outlook
To summarize, we have explicitly computed the expressions for energy values of the subgap ABSs in Josephson junctions. built with 2d and 3d triple-point fermions, in the thin-barrier limit. We have assumed a weak and homogeneous s-wave pairing in each superconducting region, which can be created via proximity effect by placing a superconducting electrode near it [41]. The barrier region can be implemented by applying a voltage of magnitude $V_{0}$ across a piece of semimetal in its normal state. By constructing the appropriate BdG Hamiltonian, we have determined the wavefunction in a piecewise continuous manner, which localizes near the S-B and B-S junctions. On enforcing the consistency of the equations obtained from matching the boundary conditions, which results from the vanishing of the relevant determinant, we have found an equation consisting of the trigonometric functions $\cos(2\beta)$ and $\sin(2\beta)$. In fact, this equation is the same for both the 2d and 3d cases. The solutions for $\cos\beta$ give the absolute energy values $|\varepsilon|$ of the ABSs. We have shown the closed-form analytical expressions for $|\varepsilon|$, which turns out to be two distinct functions [cf. Eq. (28)].
@@
We would like to point out that, although each of the two $|\varepsilon|$-expressions is a function of $\cos\phi_{12}$, the dependence is not of the form $\epsilon_{abs}^{lin}\equiv\Delta_{0}\,\sqrt{1-T_{N}\,\sin^{2}\left(\varphi_{12}/2\right)}$, found for two-band semimetals when the propagation direction is along a liearly dispersing direction [34, 36, 38] (as discussed in the introduction section). This follows from the fact that increasing the number of bands changes equations to be solved in the variables $\sin(2\beta)$ and $\cos(2\beta)$, arising as trigonomatric functions of $\beta$. In other words, when the equation reduces to the simple form of ~$\tilde{\mathcal{B}}\cos(2\beta)+\tilde{\mathcal{C}}\cos\varphi_{12}=0$, solutions of the form $\epsilon_{abs}^{lin}$ emerge. However, for multifold fermions, such a simplicity is lost, giving rise to multiple solutions of $|\varepsilon|$ as another artifact (see, for example Ref. [30], where the case of an isotropic RSW semimetal with fourfold-degenerate nodal points has been studied). An analogous increase in the complexity of the equation is also caused when the propagation direction is along a momentum direction along which the semimetal dispersion exhibit a nonlinear dependence (as seen in Ref. [40], where the case of a semi-Dirac semimetal has been investigated).
@@
In the future, it will be worthwhile to investigate the Josephson effects in twofold and multifold semimetals with nonlinear-dispersion directions [38, 39, 48, 49, 29, 50, 51], by considering the propagation of the quasiparticles/quasiholes along such an axis. Other interesting avenues to explore will be to consider S-N-S or S-B-S junctions for higher angular momentum channels (e.g., d-wave symmetric pairing channel [25]) and for FFLO pairings [37, 39] .
V Acknowledgments

### v2

@@ II.2 Results
The thin-barrier limit is physically equivalent to a Dirac-delta potential barrier. However, since we do not have any constraint on the derivatives of the componets of the wavefunction across the junctions, which follows from the linear-in-momentum dispersion [19, 31, 22, 23], the standard delta-function-potential approximation [46, 50, 43] for thin barriers cannot be used here [39, 33]. Instead, we need to start with Eq. (II.1), and impose the approximations
@@
Figure 3: Behaviour of $|\varepsilon|$ against $V_{0}$ [subfigure (a)] and $L$ [subfigure (b)], for some representative values of the remaining parameters (shown in the plotlabels).
From the four combinations of the components of the wavefunction appearing at each of the two boundaries, as shown in Eq. (II.1), we get $2\times 4=8$ linear homogeneous equations in the $8$ variables $\left\{a_{l},\,b_{l},\,a,\,b,\,c,\,d,\,a_{r},\,b_{r}\right\}$, which constitute the $8$ unknown coefficients of the wavefunction. The overall $x$-independent factors of $e^{i\,k_{y}\,y}$ cancel out equation by equation. Let $M$ be the $8\times 8$ matrix constructed from the coefficients of these 8 variables, whose explicit form is given by
@@
The consistency of the 8 equations is ensured by the condition $\text{det}\,M=0$. After some row and column operations on the matrix $M$, we can bring it to the form of $\text{det}\,\tilde{M}=0$, where
@@
$\mathcal{B}=\sin^{2}\chi\left(\cos^{4}\theta_{n}+\cos^{4}\theta_{r}\right)-\left[3\cos(2\chi)+1\right]\cos^{2}\theta_{n}\cos^{2}\theta_{r}\,,$
@@
In Fig. 2, we show the variation of $|\varepsilon|$ as a function of (I) $k_{\perp}$ (with $\varphi_{12}$ held fixed) and (II) $\varphi_{12}$ (with $k_{\perp}$ held fixed), for some representative values of $V_{0}$ and $L$. Since $\mathcal{C}$ is a function of $\cos\varphi_{12}$ [cf. Eq. (II.2)], it follows from Eq. (30) that the subgap energies are periodic in $\varphi_{12}$ with period $2\pi$, which is also evident from the curves shown in the lower panel of Fig. 2. In Fig. 3, we plot the dependence of $|\varepsilon|$ as a function of (I) $V_{0}$ (with $L$ held fixed) and (II) $L$ (with $V_{0}$ held fixed), for some representative values of $k_{\perp}$ and $\varphi_{12}$. The curves show an oscillatory behaviour, with respect to both $V_{0}$ and $L$. This is not surprising, given that the solutions are functions of $\cos(2\chi)$. Lastly,, we have plotted the behaviour of ±$\pm|\varepsilon|$ as functions of $\varphi_{12}$ and $k_{\perp}$ in the upper panel of Fig. 4, thus illustrating the dependence of the subgap energies on both these variables in a combined way. Subfigures Fig. 4(a) and Fig. 4(b) show the behaviour for two sets of values of the remaining junction variables $V_{0}$ and $L$, with $E_{F}$ kept fixed at unity.
@@
where $\epsilon_{n}$ labels the energy values of the four ABSs, arising from the two distinct values of $|\varepsilon|$. Here, $f(\zeta)=1/\left(1+e^{\frac{\zeta}{k_{B}\,T}}\right)$ is the Fermi-Dirac distribution function, with $k_{B}$ being the Boltzmann constant. The lower panel of Fig. 4 demonstrates the behaviour of $I_{J}$ as a function of $\varphi_{12}$, scaled by the appropriate numbers/variables (which we label as $\tilde{I}$), for two sets of representative parameters.
@@
The 3d pseudospin-1 quasiparticles are the generalizations of the pseudospin-1/2 quasiparticles in Weyl semimetals, with the nodal points acting as sources/sinks of the Berry flux, acting as Berry-curvature monopoles of magnitude 2 [3, 6]. They can be realized in 3d tight-binding models for cold fermionic atoms in cubic optical lattices [12] and, via ab initio simulations, have been identified in materials like TaN, NbN, and WC-type ZrTe [6, 7, 16, 8]. Analogous to Eq. (1), the effective low-energy continuum Hamiltonian, in the vicinity of a nodal point, is given by
@@
The 3d probability-current density operator takes the form of $\check{\mathbf{j}}_{3d}(\mathbf{k})=\nabla_{\mathbf{k}}\mathcal{H}_{3d}(\mathbf{k})=\mathcal{S}_{x}\,\hat{\mathbf{x}}+\mathcal{S}_{y}\,\hat{\mathbf{y}}+\mathcal{S}_{z}\,\hat{\mathbf{z}}$. The current density vector from the wavefunction $\psi\equiv\left(c_{1}\quad c_{2}\quad c_{3}\right)^{T}$ is
@@
This again leads to a vanishing contribution from the flat-band wavefunction. Eq. (34) also tells us that if the quasiparticles encounter a perpendicular barrier, say at $z=z_{0}$, while propagating along the $z$-direction, the conservation of the probability-current density leads to the two boundary conditions [23],
$\lim_{\delta\rightarrow 0}c_{1}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{1}(x,y,z_{0}+\delta)\text{ and }\lim_{\delta\rightarrow 0}c_{3}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{3}(x,y,z_{0}+\delta)\,.$ (35)
In this case, if the Fermi energy cuts the dispersion profile at an energy value $E$, the wavefunction will have the exponential factor $e^{i\,\text{sgn}(E)\,k_{z}\,z}$, where $k_{z}=\sqrt{E^{2}-k_{\perp}^{2}}$ and $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$.
III.1 S-B-S junction’s interfaces oriented perpendicular to the $z$-direction
Here we consider the propagation of the quasiparticles/quasiholes along the $z$-axis. We would like to point out that the final results would be the same if we had taken the propagation direction to be along the $x$- or $y$-axis, which follows from the isotropic nature of the dispersion. In other words, we can freely rotate our coordinate system without any physically observable consequences. The reason for our particular orientation of the coordinate axes is that the expressions for the wavefunctions and the resulting intermediate equations (to be solved) turn out to be simpler for making this choice.
Setting up the S-B-S configuration as shown in Fig. 1(b), we model the superconducting pair potential as
@@ III.1 S-B-S junction’s interfaces oriented perpendicular to the $z$-direction
where the subscripts $\{1,2,3\}$ on the fermionic creation and annihilation operators represent the three distinct band indices. Analogous to the 3d case, we assume the energy-scale hierarchies of $V_{0}\gg E_{F}\gg\Delta_{0}$ and $(V_{0}-E_{F})\gg E_{F}$.
@@
For considering normal-metal–superconductor interfaces, the $z$-component of the probability-current density operator must be modified to
$\check{\mathbf{j}}^{N}_{z}=\partial_{k_{z}}\mathcal{H}_{\text{BdG}}(\mathbf{k},\mathbf{r})=\mathcal{S}_{z}\otimes{\mathcal{S}}_{0}-{\mathcal{S}}_{0}\otimes\mathcal{S}_{z}\,.$ (40)
Analogous to the 2d case, we here infer that the electron- and hole-wavefunctions, corresponding to the flat-band state, drop out from the equations arising out of the boundary conditions. In addition, Eq. (40) translates into the following constraints, useful for relating the components from a piecewise-defined wavefunction, $\psi_{\text{BdG}}\equiv\left(c_{1}\quad c_{2}\quad c_{3}\quad c_{4}\quad c_{5}\quad c_{6}\right)^{T}$, across an interface at $z=z_{0}$:
$\lim_{\delta\rightarrow 0}c_{1}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{1}(x,y,z_{0}+\delta)\,,\quad\lim_{\delta\rightarrow 0}c_{3}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{3}(x,y,z_{0}+\delta)\,,$
$\lim_{\delta\rightarrow 0}c_{4}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{4}(x,y,z_{0}+\delta)\,,\quad\lim_{\delta\rightarrow 0}c_{6}(x,y,z_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{6}(x,y,z_{0}+\delta)\,.$ (41)
Here, we have the two boundaries at $z_{0}=0$ and $z_{0}=L$.
@@
expressed in a piecewise manner for the three regions, where we set the Fermi energy at $E_{F}$ for the corresponding normal states (i.e., for $\Delta_{0}=0$) in the regions I and III.
@@
Since the final results depend on the phase difference $\varphi_{12}\equiv\varphi_{2}-\varphi_{1}$, for simplification of the notations, we set $\varphi_{1}=0$ and $\varphi_{2}=\varphi_{12}$. Imposing the continuity of the probability current density at the junctions located at $z=0$ and $z=L$, we get the following conditions [cf. Eq. (III.1)]:
@@ III.2 Results
In order to implement the thin-barrier limit, we start with Eq. (III.1), and impose the approximations
@@
From the four combinations of the components of the wavefunction appearing at each of the two boundaries, as shown in Eq. (III.1), we get $2\times 4=8$ linear homogeneous equations in the $8$ variables $\left\{a_{l},\,b_{l},\,a,\,b,\,c,\,d,\,a_{r},\,b_{r}\right\}$, which constitute the $8$ unknown coefficients of the wavefunction. In the resulting equations, both the overall $z$-independent factors of $e^{i\left(k_{x}\,x\,+\,k_{y}\,y\right)}$, and the phase factors introduced by the action of $e^{-i\,\phi\,{\mathcal{S}}_{0}\otimes{\mathcal{S}}_{z}}$, cancel out component by component. Let $\check{M}$ be the $8\times 8$ matrix constructed from the coefficients of these 8 variables, whose explicit form is given by
@@
The consistency of the 8 equations is ensured by the condition $\text{det}\,\check{M}=0$. After some row and column operations on the matrix $\check{M}$, we can bring the constraint equation to a simple form which turns out to be exactly the same as Eq. (28). Hence, the results for the 2d case ABSs are valid for the 3d case as well. This is similar to the scenario for the pseudospin-1/2 quasiparticles/quasiholes for graphene and Weyl semimetal, where the ABS-energy expressions are identical (as we can see from Eq. (20) of Ref. [39] and Eq. (13) of Ref. [41]). Such outcomes are not surprising because, after all, both the 2d and 3d versions harbour excitations carrying the same pseudospin-values, with the continuum-limit dispersion having the same isotropic linear-in-momentum character.
The only difference from the 2d case is that, because the transverse part of the semimetallic slab is now two-dimensional (with a cross-section area of $W^{2}$), the expression for the Josephson current density across the weak link is now given by [46, 37]
[…]
where the integration is now over a 2d momentum space (analogous to the cases of Refs. [41, 33]).
@@ IV Summary and outlook
To summarize, we have explicitly computed the expressions for energy values of the subgap ABSs in Josephson junctions, built with 2d and 3d triple-point fermions, in the thin-barrier limit. We have assumed a weak and homogeneous s-wave pairing in each superconducting region, which can be created via proximity effect by placing a superconducting electrode near it [44]. The barrier region can be implemented by applying a voltage of magnitude $V_{0}$ across a piece of semimetal in its normal state. By constructing the appropriate BdG Hamiltonian, we have determined the wavefunction in a piecewise continuous manner, which localizes near the S-B and B-S junctions. On enforcing the consistency of the equations obtained from matching the boundary conditions, which results from the vanishing of the relevant determinant, we have found an equation consisting of the trigonometric functions $\cos(2\beta)$ and $\sin(2\beta)$. In fact, this equation is the same for both the 2d and 3d cases. The solutions for $\cos\beta$ give the absolute energy values $|\varepsilon|$ of the ABSs. We have shown the closed-form analytical expressions for $|\varepsilon|$, which turn out to be two distinct functions [cf. Eq. (30)].
@@
We would like to point out that, although each of the two $|\varepsilon|$-expressions is a function of $\cos\varphi_{12}$, the dependence is not of the form $\epsilon_{abs}^{lin}\equiv\Delta_{0}\,\sqrt{1-T_{N}\,\sin^{2}\left(\varphi_{12}/2\right)}$, found for two-band semimetals when the propagation direction is along a liearly dispersing direction [37, 39, 41] (as discussed in the introduction section). This follows from the fact that increasing the number of bands changes the equations to be solved in the variables $\sin(2\beta)$ and $\cos(2\beta)$, arising as trigonomatric functions of $\beta$. More explicitly, when the equation reduces to the simple form of ~$\tilde{\mathcal{B}}\cos(2\beta)+\tilde{\mathcal{C}}\cos\varphi_{12}=0$, solutions of the form $\epsilon_{abs}^{lin}$ emerge. However, for multifold fermions, such a simplicity is lost, giving rise to multiple solutions of $|\varepsilon|$ as another artifact (see, for example Ref. [33], where the case of an isotropic RSW semimetal with fourfold-degenerate nodal points has been studied). An analogous increase in the complexity of the equation is also caused when the propagation direction is along a momentum direction, along which the semimetal dispersion possesses a nonlinear dependence (as seen in Ref. [43], where the case of a semi-Dirac semimetal has been investigated).
If the barrier has impurities, structural defects, or non-uniformity, it will warrant incorporating that in the computations, which can be realistically achieved by resorting to numerics. In some simple cases, e.g., a single localized defect in the normal-phase region for graphene (cf. Ref. [38]), it is possible to derive analytical expressions for the ABS-energy levels when the radius of the defect is much much smaller than $L$. here, it was clearly seen that the Andreev levels are modified by the impurity. For our three-band semimetal, it may or may not be possible to do so, because it all depends whether we can obtain approximate forms of the resulting wavefunctions and, in the end, whether we can find the solutions from the determinant derived from the matching boundary conditions.
[…]
@@
In the future, it will be worthwhile to investigate the Josephson effects in twofold and multifold semimetals with nonlinear-dispersion directions [41, 42, 57, 58, 32, 59, 60], by considering the propagation of the quasiparticles/quasiholes along such an axis. A nontrivial topology can be exhibited when we consider an order parameter comprising multiple components that transform under the same irreducible representation, but not for the case of a single order parameter corresponding to s-wave pairing. This has been discussed in a recent preprint [61]. The topology of the gap function is manifested through its Berry-phase structure, which will have its fingerprints in the Josephson current. In the future, we would like to investigate this aspect as well. Other interesting avenues to explore will be to consider S-N-S or S-B-S junctions comprising higher-angular-momentum pairing channels (e.g., d-wave symmetric pairing channel [28]) and/or for FFLO order parameters [40, 42].
Acknowledgments